# Exploring the Limits of Fine-grained LLM-based Physics Inference via Premise Removal Interventions

### v1

@@ Abstract
Language models can hallucinate when performing complex and detailed mathematical reasoning. Physics provides a rich domain for assessing mathematical reasoning capabilities where physical context imbues the use of symbols which needs to satisfy complex semantics (e.g., units, tensorial order), leading to instances where inference may be algebraically coherent, yet unphysical. In this work, we assess the ability of Language Models (LMs) to perform fine-grained mathematical and physical reasoning using a curated dataset encompassing multiple notations and Physics subdomains. We improve zero-shot scores using synthetic in-context examples, and demonstrate non-linear degradation of derivation quality with perturbation strength via the progressive omission of supporting premises. We find that the models’ mathematical reasoning is not physics-informed in this setting, where physical context is predominantly ignored in favour of reverse-engineering solutions.
Exploring the Limits of Fine-grained LLM-based Physics Inference via Premise Removal Interventions
Jordan Meadows1, Tamsin James2, André Freitas1,3 1University of Manchester, United Kingdom 2University of Birmingham, United Kingdom 3Idiap Research Institute, Switzerland jordan.meadows@postgrad.manchester.ac.uk txj287@student.bham.ac.uk andre.freitas@idiap.ch
@@ 1 Introduction
Language models demonstrate some level of mathematical ability Lewkowycz et al. (2022); Azerbayev et al. (2023); Pan et al. (2024). This reasoning modality requires controlled symbolic behaviour involving the repeated application of mathematical operations Valentino et al. (2023); Meadows et al. (2023a), and LMs struggle to deliver this reliably Frieder et al. (2023); Liu and Yao (2024). A particularly challenging mathematical domain is that of Physics, where equation derivations serve as a rich environment within which the mathematical inference capabilities of LMs may be thoroughly examined, yet in contrast to other forms of mathematical reasoning Shakarian et al. (2023); Yuan et al. (2023); Wang and Lu (2023), there are few examples of such efforts Lewkowycz et al. (2022). While Mathematics proofs are closer to a logical argument over more abstract domain types, Physics derivations are centered around the integration of the abstraction of physical properties and laws into approximations and premises, where algebraic and calculus-related symbolic manipulation are performed to obtain novel equations, through a step-wise derivation, in close dialogue with empirical evidence.
@@
For instance, in Fig. 1, GPT-4 fails a derivation (from the present data) related to the Uncertainty Principle in Quantum Mechanics. The first error arises from attempting to substitute the RHS of the second equation into the RHS of the first, but GPT erroneously keeps the in the LHS. The second error involves an incorrect evaluation of an integral over a Dirac delta function (by a factor of ), which is a critical result in Physics. Real-world mathematical research inherently involves equation manipulation which is out-of-distribution with respect to models’ training data, either through a derivation’s specific use of notation or its underlying reasoning. If state-of-the-art LMs fail at such basic manipulation, and if current evaluation metrics fail to account for such fine-grained errors Welleck et al. (2022a); Meadows et al. (2023a), to what extent are leading methods appropriate for inference in mathematical domains Davis (2024)?
@@
(1.) A manually curated step-wise granular dataset comprising 1200 derivation steps over 218 fine-grained Physics derivations. The comprised derivations are at approximately graduate-level difficulty, and span a diverse set of subdomains including electromagnetism, classical, quantum, and statistical mechanics. Derivations are aligned to reference derivations in Wikipedia, but have been manually augmented to elicit finer steps, resulting in a highly granular reasoning corpus. The dataset qualifies premise selection steps, intermediate steps, and goal equations. As derivations are related, many can be chained together to form much longer sequences.
@@ 2 Related work
Our focus is on the generation of step-wise detailed derivations of Physics equations with LMs Brown et al. (2020); Ahmed and Devanbu (2022); Song et al. (2022); Ge et al. (2023); Hu et al. (2023); Yang et al. (2023). While we presently consider solely equations, math generation exists in various forms, split between approaches that consider formal languages First et al. (2023); Polu et al. (2022); Jiang et al. (2021); Polu and Sutskever (2020) and those considering informal mathematical natural language Ferreira and Freitas (2020); Welleck et al. (2021); Ferreira et al. (2022); Valentino et al. (2022a); Meadows and Freitas (2023). Informal reasoning approaches generally involve code generation as input to symbolic solvers He-Yueya et al. (2023); Chen et al. (2022); Drori et al. (2022); Mandlecha et al. (2022); Hu and Yu (2022); Chen et al. (2021), or directly generating math reasoning in natural language Lewkowycz et al. (2022); Welleck et al. (2022a); Chowdhery et al. (2022); Lample and Charton (2019). Numerous approaches exist for evaluating the mathematical and symbolic capabilities and robustness of models Welleck et al. (2022b); Stolfo et al. (2022); Meadows et al. (2023b), in contexts such as solving math word problems Roy et al. (2015); Liang et al. (2022); Yao et al. (2023). Various datasets exist containing mathematical reasoning Ferreira and Freitas (2020); Hendrycks et al. (2021); Welleck et al. (2021); Mishra et al. (2022) and to a lesser extent, Physics reasoning Hendrycks et al. (2021); Lewkowycz et al. (2022); Meadows et al. (2022); Pan et al. (2024). We narrow our scope to isolate equational content from natural language descriptions, for the purpose of testing purely the mechanical and mathematical capabilities of generative models in this context.
To summarise, we consider an equation Derivation Generation task, contribute a fine-grained Physics dataset spanning multiple subdomains, perturb input prompts with a Premise Removal intervention, and explore model performance and degradation due to perturbations. We aim for this data to improve the detailed step-wise equation derivation capabilities of LMs, and use it to analyse their ability to perform Physics reasoning with complex equational forms (Fig. 1 and 2). We highlight the difference between coherent mathematical and physical reasoning (Fig. 4), and discuss underlying generation degradation laws and how language models perform mathematics in this context.
3 Constructing a dataset of fine-grained Physics derivations
@@
Fig. 2 gives an example of this rewriting and the departure from natural language. With respect to granularity differences, the single intermediate step (denoted by red) in the Wikipedia derivation on the left, is actually composed of multiple fine-grained steps naturally omitted for the sake of succinct communication online. Experienced physicists may indeed skip these steps within their own workings, but there is no guarantee that language models can perform them reliably without hallucination Welleck et al. (2022a); Frieder et al. (2023); Meadows et al. (2023a); Liu and Yao (2024). On the right-hand side, this single step is expanded into four finer-grained steps that improve explainability. The first and second equations in the expanded derivation are premises extracted from the text. The third equation is formed by substituting the second premise into the first, and the fourth equation is not explicitly written anywhere – it is the Fourier transform of described non-mathematically within the initial description, which we include explicitly for completeness. The remaining equations are obtained through substitution.
We algorithmically describe this expansion and annotation process for converting online derivations into examples rendered in Appendices B-D, in Alg. 1. This annotation protocol was applied by each annotator within a double swap, review and refine setting. Initial derivations were split between each annotator, and after initial annotation, the datasets were swapped for review and changes were tracked. A second swap was performed for accepting the proposed changes.
@@
High-quality Villalobos et al. (2022) data of this kind is necessary to bridge an incompleteness gap between the communicative reasoning available online (used to train/evaluate models) and the reality of Physics calculations Meadows and Freitas (2021); Mann et al. (2018); Hopfield (1958). Establishing that models can robustly perform such reasoning is a prerequisite enabling their utility and application in theoretical discovery.
@@
We expand 60 random derivations from Wikipedia examples following the discussed protocol, and extract 218 shorter examples resembling the right-hand side of Fig. 2. These are split into three categories: Electromagnetism (containing vector calculus), Quantum Mechanics (containing Dirac notation and commutation relations), and Other Physics (containing results from classical and statistical mechanics). Tab. 1 describes the number of examples within each subdomain, and example derivations are rendered in Appendices B (electromagnetism), C (quantum), and D (other).
@@
Smaller models, such as BERT Devlin et al. (2018) and T5 Raffel et al. (2020), are limited to input sequences of up to 512 tokens. This corresponds to a LaTeX derivation comprising between 4-9 equations (excluding natural language). The original expanded derivations are segmented into shorter examples to accommodate these limitations. The relevant length distributions are given in Fig. 3 alongside that of the synthetic dataset used to fine-tune MathT5 in related work Meadows et al. (2023a), and to provide in-context examples for few-shot prompts in these experiments. Other similarities with the synthetic dataset include an overlap of 155 symbols and a similar step granularity.
@@
Premises are another crucial element of derivations. These are axiomatic equations deemed necessary for deriving a goal equation. The distribution of the number of premises per derivation is given in Fig. 4. A significant proportion of steps involves simply writing the premises and goal equation in the correct order with correct syntax (a non-trivial task for LMs Chen et al. (2024); Meadows et al. (2023a)), or substituting premises.
A final noteworthy property of the dataset is its compositionality towards longer derivations. Premises are separated from non-premises, and goal equations from some examples are premises in another, meaning that sequences may be chained together to form much longer examples of up to 20 steps (far beyond the capabilities of the state-of-the-art). The dataset itself contains both prompts and target derivations. Output text takes the form of LaTeX equations conjoined by an “and” token, which provides a minimalistic template for defining equation boundaries. The dataset is available online111https://github.com/jmeadows17/transformers-for-physics, and includes lists of derivation-specific premises and references to original Wikipedia names alongside each example.
@@ 4 Derivation Generation and Generalisation Capabilities
The experimental analysis occurs in two parts: model evaluation on an in-distribution test set in the setting of a Derivation Generation task, followed by the exploration of model performance on out-of-distribution examples obtained from the progressive removal of premises from in-distribution prompts. Perturbed prompts are out-of-distribution with respect to either a model’s training data or in-context examples. We now outline the Derivation Generation task and Premise Removal perturbations that are central to each stage of the analysis.
@@ 5.1 Derivation Generation
To supplement previous analysis, we manually evaluate the coherence of the top scoring model’s derivations in each subdomain (Tab. 4), and describe instances where reasoning failures violate laws of Physics without necessarily including mathematical errors (Fig. 5). Notably, the misuse of minus signs frequently contributes to incoherent reasoning, and equations can be recalled incorrectly even if they are given in the prompt (many are rearranged).
@@
The coherence of a derivation does not depend on the correct application of Mathematics alone. Fundamental physical assumptions guide what mathematical steps are allowed, and in many cases, such as those in Fig. 5, GPT-4 fails to understand this. For example, the Electromagnetism excerpt is only valid if the electric field and charge density are constant throughout space. In this case, all integrals indeed cancel to give Gauss’ law. However, as this general law is true when quantities are spatially dependent, this reasoning is flawed. In the Quantum excerpt, both equations are true simultaneously only if and commute. Assuming each are respectively the quantum mechanical operators representing position and momentum, GPT-4 enforcing this commutativity violates Heisenberg’s uncertainty principle. The Classical excerpt is from a derivation that attempts to obtain Snell’s law. GPT-4 keenly asserts that light cruises across optical boundaries at 1 m/s without refraction.
@@ 5.2 Premise Removal
To give examples, in the supplementary material222https://github.com/jmeadows17/transformers-for-physics/blob/main/Supplementary_Model_Derivations.pdf, the key concept in C.89 is to equate kinetic energy with work and rearrange for velocity, but rather than introduce kinetic energy as a premise, GPT reverse-engineers the derivation from the goal equation, which we mark as correct. The derivations A.54 and B.54 would be marked as correct if they did not skip the premise , but the reasoning as it stands would allow , which is false.
Overall, the manual evaluation gives a liberal estimate of model accuracy on the dataset in Tab. 6. We would expect Physics professionals to give lower scores, particularly for .
The non-linear degradation in accuracy due to the gradual removal of premises aligns closely with scores obtained from the automatic metrics (notably BLEU). This supports our assumption in Section 4.1, and suggests machine translation and text summarisation metrics may serve some use in this context, despite their obvious flaws when evaluating individual examples and comparing models’ math capabilities.
@@
Working backwards from the result, this reasoning approach is typically characterised by a lack of understanding of the problem. Due to the prevalence of this reasoning in the output this is perhaps the core mechanism behind how language models solve this task. Such a solution is given below:
@@ 6 Conclusion
We explore the mathematical ability of language models in the context of a derivation generation task, using a novel dataset of equation derivations in the Physics domain. The difference between steps is small, only one or two operations, with the intent of emulating granular mathematical workings and exploring how models perform detailed calculations involving a variety of notations specific to fields such as electromagnetism and quantum mechanics.
We few-shot prompt GPT models with synthetic in-context examples which increases their scores obtained from text generation metrics, and manually evaluate hundreds of generated derivations (in the context of a premise removal intervention on the prompt) revealing a non-linear relationship between the perturbation strength and average derivation quality, as reported by manual and automatic scores. However, manual and automatic scores disagree on the relative difficulty of Physics subdomains (as represented in the dataset).
@@
We find that language models (particularly GPT-4) derive goal equations from premises through reverse-engineering intermediate steps without appropriate consideration of basic underlying Physics. This becomes increasingly apparent as we progressively remove premises from few-shot prompts (without perturbing premises from in-context examples).
@@
In particular, many algebraic errors arise from attempts at substitution. While it is certainly a non-trivial task building synthetic datasets that reflect all aspects of mathematical reasoning Toshniwal et al. (2024), it is more manageable to focus on the application of individual operations (such as substitution), and use symbolic engines to apply them in the context of vast vocabularies of symbols and notations representative of target subdomains, which could potentially bolster the out-of-distribution mathematical abilities of transformer-based models.
@@
Although exploration of the synergy between language models and symbolic engines is underway Yang et al. (2024); Davis (2024), Physics derivations rely on assumptions of physical systems which are either implicit (e.g., commutative properties of quantum operators) or stated explicitly in prompts (e.g., a particular premise equation). These assumptions govern the reasoning paths allowed in derivations, and are hence independent of the ability to perform fine-grained algebraic manipulation. This is effectively a premise/operation selection problem, and integrating symbolic solvers does not inherently improve this limitation Liu and Yao (2024). We find contemporary language models severely lacking in this regard, and suggest research efforts should be directed towards LM-based querying of appropriate knowledge bases during inference.

### v2

@@ Abstract
Language models (LMs) can hallucinate when performing complex mathematical reasoning. Physics provides a rich domain for assessing their mathematical capabilities, where physical context requires that any symbolic manipulation satisfies complex semantics (e.g., units, tensorial order). In this work, we systematically remove crucial context from prompts to force instances where model inference may be algebraically coherent, yet unphysical. We assess LM capabilities in this domain using a curated dataset encompassing multiple notations and Physics subdomains. Further, we improve zero-shot scores using synthetic in-context examples, and demonstrate non-linear degradation of derivation quality with perturbation strength via the progressive omission of supporting premises. We find that the models’ mathematical reasoning is not physics-informed in this setting, where physical context is predominantly ignored in favour of reverse-engineering solutions.
@@ 1 Introduction
Language models demonstrate some level of mathematical ability Lewkowycz et al. (2022); Liu et al. (2023); Azerbayev et al. (2023); Pan et al. (2024). This reasoning modality requires controlled symbolic behaviour involving the repeated application of mathematical operations Valentino et al. (2023); Meadows et al. (2023a), and LMs struggle to deliver this reliably Frieder et al. (2023); Liu and Yao (2024). A particularly challenging mathematical domain is that of Physics, where equation derivations serve as a rich environment within which the mathematical inference capabilities of LMs may be thoroughly examined, yet in contrast to other forms of mathematical reasoning Shakarian et al. (2023); Yuan et al. (2023); Wang and Lu (2023), there are few examples of such efforts Lewkowycz et al. (2022). While Mathematics proofs are closer to a logical argument over more abstract domain types, Physics derivations are centered around the integration of the abstraction of physical properties and laws into approximations and premises, where algebraic and calculus-related symbolic manipulations are performed to obtain novel equations, through a step-wise derivation, in close dialogue with empirical evidence.
@@
For instance, in Fig. 1, GPT-4 fails a derivation (from the present data) related to the Uncertainty Principle in Quantum Mechanics. The first error arises from attempting to substitute the RHS of the second equation into the RHS of the first, but GPT erroneously keeps the in the LHS. The second error involves an incorrect evaluation of an integral over a Dirac delta function (by a factor of ), which is a critical result in Physics. Real-world research inherently involves equation manipulation which is out-of-distribution with respect to models’ training data, either through a derivation’s specific use of notation or its underlying reasoning. If state-of-the-art LMs fail at such basic manipulation, and if current evaluation metrics fail to account for such fine-grained errors Welleck et al. (2022a); Meadows et al. (2023a), to what extent are leading methods appropriate for inference in mathematical domains Davis (2024)?
@@
(1.) A manually curated step-wise granular dataset comprising 1200 derivation steps over 218 fine-grained Physics derivations at approximately graduate-level difficulty. Examples span a diverse set of subdomains including electromagnetism, classical, quantum, and statistical mechanics. The derivations are aligned to reference those on Wikipedia, but have been manually augmented to include finer steps, and are mapped to prompts containing premises and goal equations.
@@ 2 Related work
Our focus is on the generation of step-wise detailed derivations of Physics equations with LMs Brown et al. (2020); Ahmed and Devanbu (2022); Song et al. (2022); Ge et al. (2023); Hu et al. (2023); Yang et al. (2023). While we presently consider solely equations, math generation exists in various forms, split between approaches that consider formal languages First et al. (2023); Polu et al. (2022); Jiang et al. (2021); Polu and Sutskever (2020) and those considering informal mathematical natural language Ferreira and Freitas (2020); Welleck et al. (2021); Ferreira et al. (2022); Valentino et al. (2022a); Meadows and Freitas (2023). Informal reasoning approaches generally involve code generation as input to symbolic solvers He-Yueya et al. (2023); Chen et al. (2022); Drori et al. (2022); Mandlecha et al. (2022); Hu and Yu (2022); Chen et al. (2021), or directly generating math reasoning in natural language Lewkowycz et al. (2022); Welleck et al. (2022a); Chowdhery et al. (2022); Lample and Charton (2019). Numerous approaches exist for evaluating the mathematical and symbolic capabilities and robustness of models Welleck et al. (2022b); Stolfo et al. (2022); Meadows et al. (2023b), in contexts such as solving math word problems Roy et al. (2015); Liang et al. (2022); Yao et al. (2023). Various datasets exist containing mathematical reasoning Ferreira and Freitas (2020); Hendrycks et al. (2021); Welleck et al. (2021); Mishra et al. (2022), and to a lesser extent, Physics Hendrycks et al. (2021); Lewkowycz et al. (2022); Meadows et al. (2022); Pan et al. (2024). We narrow our scope to isolate equational content from natural language descriptions, for the purpose of testing purely the equation manipulation capabilities of LMs in the Physics domain.
To summarise, we consider an equation Derivation Generation task, contribute a fine-grained Physics dataset spanning multiple subdomains, perturb input prompts with a Premise Removal intervention, and explore model performance and degradation due to perturbations. We aim for this data to improve the detailed step-wise equation derivation capabilities of LMs, and use it to analyse their ability to perform Physics reasoning with complex equational forms (Fig. 1 and 2). We later highlight the difference between coherent mathematical and physical reasoning (Fig. 4), and discuss underlying generation degradation laws and how LMs perform mathematics in this context.
3 Physics Dataset Construction
@@
Fig. 2 gives an example of this rewriting and the departure from natural language. With respect to granularity differences, the single intermediate step (denoted by red) in the Wikipedia derivation on the left, is actually composed of multiple fine-grained steps naturally omitted for the sake of succinct communication online. Experienced physicists may indeed skip these steps within their own workings, but there is no guarantee that LMs can reliably perform them without error Welleck et al. (2022a); Frieder et al. (2023); Meadows et al. (2023a); Liu and Yao (2024); Quan et al. (2024). On the right-hand side, this single step is expanded into four finer-grained steps that improve explainability. The first and second equations in the expanded derivation are premises extracted from the text. The third equation is formed by substituting the second premise into the first, and the fourth equation is not explicitly written anywhere – it is the Fourier transform of described non-mathematically within the initial description now included for completeness. The remaining equations are obtained through substitution.
We algorithmically describe this expansion and annotation process for converting online derivations into examples rendered in Appendices B-D, in Alg. 1. This protocol was applied by each annotator within a double swap, review and refine setting. Initial derivations were split between each annotator, and after initial annotation, the datasets were swapped for review and changes were tracked. A second swap was performed for accepting the proposed changes.
@@
High-quality Villalobos et al. (2022) data of this kind is necessary to bridge an incompleteness gap between the communicative reasoning available online (used to train/evaluate models) and the reality of Physics calculations Pan et al. (2024); Akrobotu et al. (2022); Meadows and Freitas (2021); Mann et al. (2018); Hopfield (1958). Establishing that models can robustly perform such reasoning is a prerequisite enabling their utility and application in theoretical discovery.
@@
We expand 60 derivations from Wikipedia examples following the discussed protocol, and extract 218 shorter examples resembling the right-hand side of Fig. 2. These are split into three categories: Electromagnetism (containing vector calculus), Quantum Mechanics (containing Dirac notation and commutation relations), and Other Physics (containing results from classical and statistical mechanics). Tab. 1 describes the number of examples within each subdomain, and example derivations are rendered in Appendices B (electromagnetism), C (quantum), and D (other).
@@
Smaller models, such as BERT Devlin et al. (2018) and T5 Raffel et al. (2020), are limited to input sequences of up to 512 tokens. This corresponds to a LaTeX derivation comprising between 4-9 equations. The original expanded derivations are segmented into shorter examples to accommodate these limitations. The relevant length distributions are given in Fig. 3 alongside that of the synthetic dataset used to fine-tune MathT5 in related work Meadows et al. (2023a), and to provide in-context examples for few-shot prompts in these experiments. Other similarities with the synthetic dataset include an overlap of 155 symbols and a similar step granularity.
@@
Premises are another crucial element of derivations. These are axiomatic equations deemed necessary for deriving a goal equation. The distribution of the number of premises per derivation is given in Fig. 4. A significant proportion of steps involves simply writing the premises and goal equation in the correct order with correct syntax (a non-trivial task for LMs Chen et al. (2024); Meadows et al. (2023a)), or substituting expressions.
A final noteworthy property of the dataset is its compositionality towards longer derivations. Premises are separated from non-premises, and goal equations from some examples are premises in another, meaning that sequences may be chained together to form much longer examples of up to 20 steps. The dataset itself contains both prompts and target derivations. Output text takes the form of LaTeX equations conjoined by an “and” token, which provides a minimalistic template for defining equation boundaries. The dataset is available online111https://github.com/jmeadows17/transformers-for-physics, and includes lists of derivation-specific premises and references to original Wikipedia names alongside each example.
@@ 4 Derivation Generation and Generalisation Capabilities
The experimental analysis occurs in two parts: evaluation on an in-distribution test set in the setting of a Derivation Generation task, followed by an out-of-distribution evaluation facilitated by the progressive removal of premises from in-distribution prompts. Perturbed prompts are out-of-distribution with respect to either a model’s training data or in-context examples.
@@ 4.2 Controlled Premise Removal
The goal of premise removal is to systematically remove crucial mathematical and physical context from prompts, such that a given model must either recall this missing context and use it appropriately, or derive the goal equation via an alternative route. If a model understands the underlying reasoning, then the removal of premises should not incur derivation errors. If errors do occur as a function of the number of premises removed, we can study how certain reasoning capabilities degrade as the strength of the perturbation increases.
@@ 5.1 Derivation Generation
To supplement previous analysis, we manually evaluate the coherence of the top scoring model’s derivations in each subdomain (Tab. 4), and describe instances where reasoning failures violate laws of Physics without necessarily including mathematical errors (Fig. 5). Notably, the misuse of minus signs frequently contributes to incoherent reasoning, and equations can be recalled incorrectly even if they are given in the prompt.
@@
The coherence of a derivation does not depend on the correct application of Mathematics alone. Fundamental physical assumptions guide which mathematical steps are allowed, and in many cases, such as those in Fig. 5, GPT-4 fails to understand this. For example, the Electromagnetism excerpt is only valid if the electric field and charge density are constant throughout space. In this case, all integrals indeed cancel to give Gauss’ law. As this law is true when quantities are spatially dependent, this reasoning is flawed. In the Quantum excerpt, both equations are true simultaneously only if and commute. Assuming each are respectively the quantum mechanical operators representing position and momentum, GPT-4 violates Heisenberg’s uncertainty principle. The Classical excerpt is from a derivation that attempts to obtain Snell’s law. GPT-4 asserts that light cruises across optical boundaries at 1 m/s without refraction.
@@ 5.2 Premise Removal
To give helpful examples, in the supplementary material222https://github.com/jmeadows17/transformers-for-physics/blob/main/Supplementary_Model_Derivations.pdf, the key concept in C.89 is to equate kinetic energy with work and rearrange for velocity, but rather than introduce kinetic energy as a premise, GPT reverse-engineers the derivation from the goal equation, which we mark as correct. The derivations A.54 and B.54 would be marked as correct if they did not skip the premise , but the reasoning as it stands would allow , which is false.
Overall, the manual evaluation gives a liberal estimate of model accuracy on the dataset in Tab. 6. We would expect Physics professionals to give lower scores, particularly for . The non-linear degradation in accuracy due to the gradual removal of premises aligns closely with scores obtained from the automatic metrics (notably BLEU).
@@
Working backwards from the result, this approach is characterised by a lack of understanding of the problem. Due to the prevalence of this reasoning in the outputs, this is perhaps the core mechanism behind how language models derive given equations. An example is given below:
@@ 6 Conclusion
We explore the mathematical ability of LMs in the context of a derivation generation task using a novel dataset of equation derivations and prompt interventions. The difference between reasoning steps is only one or two operations, with the intent of emulating granular mathematical workings and exploring how models perform detailed calculations involving a variety of Physics notations.
@@
We apply premise removal interventions on prompts to reveal a non-linear relationship between the perturbation strength and average derivation quality, as reported by manual and automatic scores. We find that models (particularly GPT-4) derive goal equations from premises through reverse-engineering intermediate steps without appropriate consideration of basic underlying Physics. This becomes increasingly apparent as we progressively remove premises from few-shot prompts.
@@
Furthermore, many algebraic errors arise from attempts at substitution. While it is challenging to build synthetic datasets that reflect all aspects of mathematical reasoning Toshniwal et al. (2024), it is more manageable to focus on the application of individual operations (such as substitution), and use symbolic engines to apply them in the context of vast vocabularies of symbols and notations representative of target subdomains, which could potentially bolster models’ out-of-distribution mathematical abilities. Our use of synthetic in-context examples (involving substitution) improves models’ evaluation scores.
@@
Although exploration of the synergy between language models and symbolic engines is underway Yang et al. (2024); Davis (2024), Physics derivations rely on physical behaviour which is implicitly assumed (e.g., commutative properties of quantum operators), yet is ignored by LMs. Such assumptions govern the reasoning paths allowed in derivations, and are hence independent of the ability to perform fine-grained algebraic manipulation. This is effectively a premise/operation selection problem, and integrating symbolic solvers does not inherently improve this limitation Liu and Yao (2024). We find contemporary language models severely lacking in this regard, and suggest research efforts should be directed towards LM-based querying of appropriate knowledge bases during inference alongside solver integration Trinh et al. (2024).
7 Limitations
We have not extensively optimised prompts for model performance. However, our few-shot approach mirrors related work Azerbayev et al. (2023); Meadows et al. (2023a). Our focus on fine-grained mathematical transformations between equations excludes natural language (e.g., Fig. 2) as a design choice. Text generation metrics (e.g., ROUGE) fail to account for mathematical errors in individual derivations, however we contrast this with a human evaluation of 300 derivations in the case of GPT, which supports our conclusions. Due to the small number of human annotators responsible for curating the dataset and manually evaluating model outputs, both elements of this research may vary with the number of annotators or their experience in Physics. Our focus in this regard is a self-consistent set of derivations and assessment criteria agreed upon by the annotators involved.
Overall ethical impact.
This work explores a systematic way to elicit the mathematical/symbolic inference properties of Transformer-based models in a mathematical language processing task. As such, it contributes in the direction of a critique of the reasoning capabilities and the biases of these models, particularly in the Physics domain.